OpenStudy (anonymous):

Mrs. Jordan's car uses 6 gallons of gas to travel 150 miles how much gas will be needed a 4 500 mile trip?

OpenStudy (studygurl14):

Set up a proportion. Is that a 4,500 mile trip?

OpenStudy (tkhunny):

Dimensional Analysis!! \(\dfrac{6\;gallons}{150\;miles}\cdot\dfrac{4500\;miles}{1\;trip}\) Very clean. Very straight-forward. Learn it. Trust me on this.